OpenStudy (anonymous):

Find the length of c if angle A=62 degrees angle B=39 degrees and b = 15.

OpenStudy (anonymous):

draw a triangle so you can figure out what you are trying to find hard just with numbers and no picture

OpenStudy (anonymous):

you got one?

OpenStudy (anonymous):

drawing it :)

OpenStudy (anonymous):

ok, draw it here then we can do it

OpenStudy (anonymous):



OpenStudy (anonymous):

thats about what it looks like and this is trig that im working on aswell

OpenStudy (anonymous):



OpenStudy (anonymous):

you want to use the law of sines \[\frac{a}{\sin(A)}=\frac{b}{\sin(B)}=\frac{c}{\sin(C)}\] you know \(A, b, B\) but you don't know \(C\) except you do what is \(C\) ?

OpenStudy (anonymous):

i have no idea but yes its law of sines

OpenStudy (anonymous):

of course you do

OpenStudy (anonymous):

i meant \(A=62,B=39,C=?\)

OpenStudy (anonymous):

if it is not clear, i will tell you

OpenStudy (anonymous):

79

OpenStudy (anonymous):

whew

OpenStudy (anonymous):

thats right because triangles =180 degrees :)

OpenStudy (anonymous):

so \[\frac{15}{\sin(39)}=\frac{c}{\sin(79)}\]

OpenStudy (anonymous):

you have 3 out of the 4 numbers, you can find the fourth \[c=\frac{15\sin(39)}{\sin(79)}\] and a calculator

OpenStudy (anonymous):

14.457/ -0.444

OpenStudy (anonymous):

correct?